OpenStudy (doc.brown):

Odds: I can choose either a deck of 2 or 4 cards. The cards say what chance I have of succeeding. The deck of 2 cards has a 50% and 75%. The deck of 4 has 45% 50% 75% and 80%. Which deck should I pick to increase my odds for success?

OpenStudy (doc.brown):

In the end I will have to choose only one card from the chosen deck.

OpenStudy (kmeis002):

You need to calculate the Expected value of each deck, or what value you can expect if choosing from each deck. \[EV= \sum x_ip_i \] where x= value of a card aand p = probability of obtaining that card. So for case 1 (2 card deck): Our values for x are .5 and .75, and the probability of drawing each card is 0.5 so \[EV_{deck1} = (0.5)(0.5) + (0.75)(0.5) = .625 \] We can expect a value of 0.625, or in this case a 0.625% success rate Deck 2, the values can be found by looking, but the probability decreases to 0.25: \[ EV = (0.25)(0.45) + (0.25)(.5) + (0.25)(0.75) + (0.25)(0.8) = 0.625\] So what do these results say about which deck we should choose?